OpenStudy (anonymous):

Solve: -1/2 x + 1 = −x + 8

OpenStudy (anonymous):

first add one to both sides

OpenStudy (anonymous):

-1/2x + 1 = -x + 8 -1 -1

OpenStudy (anonymous):

i meant subtract

OpenStudy (campbell_st):

start by subtracting 1 from both sides -1/2 x = -x + 7 add x to both sides 1/2 x = 7 I'll let you finish

OpenStudy (anonymous):

so \[-1/2x=-x+9?\]

OpenStudy (anonymous):

i mean -x+7